OpenStudy (anonymous):

sinx+2cos(2y)=1 implicit differentiation ?

OpenStudy (anonymous):

We have to derive everything and then solve for y'

OpenStudy (anonymous):

which is cosx (-4sin(2y))y' = 0

OpenStudy (anonymous):

And we put y' wherever we integral with respect to y.... So we now solve for y' (-4sin(2y))y' = -cosx

OpenStudy (anonymous):

y' = (1/4)(cosx/sin(2y)

OpenStudy (freckles):

What happen to add operation between sin(x) and 2cos(2y)?

OpenStudy (freckles):

(u+v)'=u'+v' not u'v'

OpenStudy (freckles):

\[\cos(x)-4\sin(2y) \cdot y'=0 \text{ try silving this for } y'\]

OpenStudy (freckles):

And you still get the same thing as what master got somehow.